How would you do a linear system that has one equation with a fraction, using the elimination method? Such as the following: 3/2x + y= -5/2
4x + y= -5?
Your help is much appreciated.

Answer by stanbon(75887)

When you have an equation you NEVER have to
deal with fractions.
-----
3/2x + y= -5/2
4x + y= -5
---------
Multiply thru by 2 to get:
3x + 2y = -5
4x + y = -5
----
Multiply thru the 2nd Eq. by 2:
3x + 2y = -5
8x + 2y = -10
---
Subtract and solve for "x":
5x = -5
x = -1
----
Solve for "y":
4x+y = -5
4(-1) + y = -5
-4 + y = -5
y = -1
----
Solution: (-1,-1)
=======================
Cheers,
Stan H.
===========